OpenStudy (anonymous):

On a Airline, Approximatley 10% Of The Airline Passengers Who Are Booked For A Flight Do Not Show Up For The Flight . The Airline has Booked 160 Passengers For A Flight With Maximum seating Of 135. How Many Of The 160 Passengers Booked For This Flight With NOT Have A Seat, Assuming 10% Of The Booked Passengers Do Not Show Up. ?

OpenStudy (anonymous):

160x(1-0.1)-135

OpenStudy (anonymous):

How'd you Get That Though ?

OpenStudy (anonymous):

10% of the booked passengers are dead and gone , so only (100%-10%/ 1-0.1)90% of the passengers will show up. out of that 90%, 135 people will have a seat.

OpenStudy (anonymous):

Ohhh But Where Did You Get The 0.1?

OpenStudy (anonymous):

10% =10/100 = 0.1 ?

OpenStudy (anonymous):

Oh Alright Sorry Math Is My Worst subject

OpenStudy (anonymous):

LOL it is all good

OpenStudy (anonymous):

when Mr.Lee Purchased His Car, It Had A Value Of $15,000. In Each Of The First Two Years After He Purchased It, Its Value decreased By 10% Of The Previous Years Value. #of Years l Value Of Car 0 15,000$ 1 13,500$ 2 12,150$ 3 ? How Do I Find Out What The Next One IS?

OpenStudy (anonymous):

first you try to at what rate is the value decreasing; 12150/ 13500 = 90% 13500/15000=90%. As you can see, after a year, the value is only 90% of that in the previous year. So next one will be 12100X90%

OpenStudy (anonymous):

But The Answers I Have Here Say A)1215 B)4500 C)12150 D)10935 And When I Did 12100x90% It Got Me 10890?

OpenStudy (anonymous):

sorry ..there is typo is should be 12150X90%

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

Ohhhhhhhh

OpenStudy (anonymous):

sorry ^ infinity haha

OpenStudy (anonymous):

Lmao It's Okay I'm Just Glad Your Helping Me Everybody Else areJerks

OpenStudy (anonymous):

haha yay i am not a jerk.. :P

OpenStudy (anonymous):

lmao math sucks though

OpenStudy (anonymous):

NO:P

OpenStudy (anonymous):

I Used To Be SO Good In Math Then I Got A Teacher That I Couldn't Understand And Strated Slacking Then I Failed -_-

OpenStudy (anonymous):

LOL you should have switched your class or seomthing